OpenStudy (tennistar):

For which value of x does the graph of y = 2x2 - 7x + 6 cross the x-axis?

OpenStudy (tennistar):

answers choices

OpenStudy (nenadmatematika):

what is the y coordinate of the point which lies on the x axis?

OpenStudy (tennistar):

0

OpenStudy (lgbasallote):

Hint!! the x-intercept is the value of x where the graph crosses the x-axis do you know how to take the x-intercept?

OpenStudy (anonymous):

\[(x-2) (2 x-3)=0 \]\[\left\{x=\frac{3}{2},x=2\right\} \]See the attached plot.